Determine a region of the xy-plane for which the given differential equation would have a unique solution whose graph passes through a point (x0, y0)in the region.


(x^2+y^2)y prime=y^2


1
Expert's answer
2021-11-03T15:12:06-0400

According to theorem of Existence of Unique Solution if "f(x,y)" and "df\/dy" are continuous

on rectangular region "R" then there is exist interval "I" on which unique exists.


We have:

"f(x,y)=\frac {y^2}{x^2+y^2}"


"\frac {df}{dy}=\frac {2y(x^2+y^2)-y^2(x^2+2y)}{(x^2+y^2)^2}"


"f(x,y)" and "df\/dy" are not continuous at (0,0)


So, a unique solution exists in the region consisting of all points in the xy-plane except (0,0)